5. A cart is at x = 5.25m and time t = 0. The cart accelerates at


4.15m/s^2. If The speed of the cart at t=0 is 3m/s, find the position


of the cart at t =2s and also determine where the cart is when it


reaches a speed of 5m/s. (3 points)


6. A stone is thrown vertically up with an initial velocity of 4.9m/s from


the top of the building that is 64m high. On its way down, it misses


the top of the building and goes straight to the ground. Find:


(9 points)


a. Its maximum height relative to the ground


b. Its time of flight or the total time it is in air, and


c. Its velocity just before it reaches the ground.

1
Expert's answer
2022-05-01T17:17:14-0400

5.


"x(t)=x_0+v_0t+\dfrac{at^2}{2}"

"v(t)=v_0+at"

a)


"x(2)=5.25+3(2)+\dfrac{4.15(2)^2}{2}=19.55(m)"

b)


"v(t)=3+4.15t=5"

"t=\dfrac{2}{4.15} s"

"x(\dfrac{2}{4.15} )=5.25+3(\dfrac{2}{4.15} )+\dfrac{4.15(\dfrac{2}{4.15} )^2}{2}"

"=\dfrac{29.7875}{4.15} (m)\approx7.178(m)"

6.

a.


"h_{max}=h_0+\dfrac{v_0^2}{2g}"

"h_{max}=64m+\dfrac{(4.9m/s)^2}{2(9.8m/s^2)}=65.225m"

b.


"h(t)=h_0+v_0t-\dfrac{gt^2}{2}"

"64+4.9t-\dfrac{9.8t^2}{2}=0"

"t^2-t-\dfrac{64}{4.9}=0, t\ge0"

"t=\dfrac{1+\sqrt{1+4(\dfrac{64}{4.9})^2}}{2}\approx13.57(s)"

c.

"|v_{ground}|=\sqrt{2gh_{max}}"

"|v_{ground}|=\sqrt{2(9.8m/s^2)(65.225m)}\approx35.755m/s"